OpenStudy (anonymous):

What are the solutions to\[x^2 =-1\]

OpenStudy (anonymous):

x = sqrt(-1) x = +i, -i

OpenStudy (kira_yamato):

x = ±i

OpenStudy (agreene):

\[x^2=-1\]\[\sqrt{x^2}=\sqrt{-1}\]\[x= i\] Note that it is NOT plus minus... i is defined as -1^(1/2) to get -i you have to do something like: \[-\sqrt{-1}\]

OpenStudy (anonymous):

but (-i)^2 = -1 so -i is a solution

OpenStudy (anonymous):

along with plain i